(a) A sphere has a surface uniformly charged with 1.00 C. At what distance from its center is the potential 5.00 MV? (b) What does your answer imply about the practical aspect of isolating such a large charge?


1
Expert's answer
2022-01-18T13:42:30-0500

(a) From the equation for potential, find the distance:


"V=\dfrac{Q}{4\pi\epsilon_0R}→R=\dfrac{Q}{4\pi\epsilon_0V}=1798\text{ m}."


(b) Isolating such a large charge requires a lot of space.